# Tidally Delayed Spin-Down of Very Low Mass Stars

### v2

@@ 2.3 Combining magnetic braking and tides
×$\times(1-2{\rm H}(\tilde{a}_{\rm crit}-\tilde{a}(t))).$ (17)
To account for the demise of the companion when it reaches the critical orbital separation $\tilde{a}_{\rm crit}$ defined as the larger of the Roche limit (Eggleton, 1983) or stellar radius, we include a multiplicative function $(1-2{\rm H}(\tilde{a}_{\rm crit}-\tilde{a}(t)))$ in the tidal torque term, where ${\rm H}(\tilde{a}_{\rm crit}-\tilde{a}(t))$ is the logistic function given by
@@
While the magnetic braking term always decreases the stellar spin, the influence of the tidal term depends on the orbital evolution of the secondary. If the secondary moves outward, then the orbital angular momentum will increase and tidal force will act to spin down the star. But if the secondary moves inward, reducing the orbital angular momentum, then tides increase the stellar spin. This evolution is thought to be much more efficient for cool stars with outer convection zones than for hot stars with outer radiative layers (Goupil and Zahn, 2008).
2.4 Fixed model parameters
To solve for the stellar spin evolution from the system of equations (1), (2), (4) (5), (7), (15) and (17) (or (3) replacing (17) when only magnetic braking is considered) there are several parameters that we keep fixed for all of the cases that we solve: (i) We take $0.009\leq\epsilon\leq 0.0036$ which is consistent with model-inferred Rossby numbers for the transition to the X-ray saturated regime of both partially and fully convective stars (Wright et al., 2011; Reiners et al., 2014; Wright et al., 2018) and the range accounts for the corresponding uncertainty in the saturation Rossby number. (ii) we take $\lambda=1/3$, consistent with the range inferred from observations of star spot covering fractions for FGKM stars (Nichols-Fleming and Blackman, 2020). (iii) We use $q=1$ for all stars, which is equivalent to assuming that the full stellar mass is coupled to the anchoring magnetic field that contributes to spin-down. (iv) For the radius of gyration $\beta_{\star}$ and apsidal motion constant $\rm k$, we use values from stellar evolutionary models, where the evolutionary tracks are computed using the Modules for Experiments in Stellar Astrophysics package (MESA) (Claret, 2023). For a stellar mass $M_{\star}=0.2M_{\odot}$ we use $\beta_{\star}=0.458$ and $\rm k=0.155$, while for $M_{\star}=0.6M_{\odot}$ we take $\beta_{\star}=0.359$ $\rm k=0.054$ (vi) For the convective turnover timescale $\rm\tau_{c}$ we use the prescription given in (Wright et al., 2018).
@@ 3.1 Description of systems studied
To study M dwarf-BD and K dwarf-Jupiter mass companion systems we numerically solve the generalized system of equations (1), (2), (4) or (5), (7), (15) and (17) derived in section 2, which describes the stellar spin evolution when spin-down due to magnetic braking is coupled to tides. To distinguish between fully and partially convective cases, we vary stellar mass and take different values for certain fixed model parameters, while solving the same set of equations. We assume no significant difference in the dynamo mechanisms operating in fully versus partially convective stars, as current observations provide no strong evidence for such a distinction. In particular, the relationship between X-ray luminosity and Rossby number shows a similar trend for both cases (Wright et al., 2018), suggesting comparable magnetic activity. Any dynamo-related differences are beyond the present scope. For all investigations below, we compare cases with both magnetic braking and tides to cases with just magnetic braking. For the latter, we solve equation (3) instead of equations (15) and (17). Our model parameters are given in section 2.4 and the initial system properties for each case are given in Table 1. Using these values, we vary the companion mass and fix the initial value of the orbital separation, or vice versa. We define the initial orbital separation $a_{\rm i}$ as a fraction $p_{\rm f}$ of the corotation radius $a_{\rm i}=p_{\rm f}R_{\rm co}$, where the corotation radius is given by $R_{\rm co}=(G(M_{\star}+m)/\Omega^{2}_{\star})^{1/3}$.
@@
$P_{\star}/\rm Days$ 0.1 0.3 6 10 (Popinchalk et al., 2021)
$B_{\rm p\star}/\rm G$ 100 50 200 5 (Kochukhov, 2020)
$\dot{M_{\star}}/\dot{M}_{\odot}$ 0.1 0.1 0.1 1 (Wood et al., 2021)
@@ 3.2 Spin evolution for fully convective stars
Figure 1: Evolution of rotation and orbital separation for System 1. We use a $0.2M_{\odot}$ mass M-dwarf with $0.1$ days initial period, $100\rm G$ magnetic field and companion BD at $a_{\rm i}=6R_{\rm co}$ initial orbital separation. Different colors represent varying BD companion masses between $15-75M_{\rm J}$. Top panel: spin evolution due only to magnetic braking (grey) and coupled to tides (colored). Cusps represent engulfment points. The thickness of the grey and colored curves is determined by the highest and lowest $\epsilon$ values (see text). Black dots represent fully convective M dwarfs with measured rotation periods and estimated ages from Galactic kinematics (Irwin et al., 2011). The vertical grey dashed lines separate M-dwarfs in the thin, intermediate and thick disk populations with corresponding age ranges of $0.5-3$ Gyr, $3-7$ Gyr and $7-13$ Gyr. The x-axis and y-axis are normalized to 1 Gyr and a 1-day rotation period, respectively. Bottom panel: orbital separation for solutions without tides (dashed) and coupled to tides (solid) analgous to the top panel. The x-axis is normalized to 1 Gyr, and the y-axis is normalized to the stellar radius $R_{\star}$.
@@
The top and bottom panels of Fig. 1 show the time evolution of the rotation period and orbital separation for system 1 of Table 1. For initial values we take a rotation period of $P_{*}=0.1$ days, a magnetic field $B_{\rm p,\star i}=100\rm G$ and orbital separation $a_{\rm i}=6R_{\rm co}\approx 0.015\rm AU$. We consider cases of companion BD masses between $15-75M_{\rm J}$, each represented by a different color. Mass accretion processes are not considered in the current study, and may be important for large companion masses. We take $75M_{\rm J}$ BD dwarf as our largest companion mass case for visualization and to serve as a conceptual bridge towards future work involving stellar binaries and mass transfer. Black dots are data points from (Irwin et al., 2011) with measured rotation periods, and estimated ages from galactic kinematics for 41 field M-dwarfs. This data set typifies the observed trend, and provides a simple comparison to our results, given the availability of both rotation periods and ages in the sample. In Fig. 1, the vertical grey dashed lines represent boundaries between Galactic thin, intermediate, and thick disk M-dwarf populations with assigned ages $0.5-3$ Gyr, $3-7$ Gyr, and $7-13$ Gyr, respectively. The data points in Fig. 1 use the average value of these disk sub-population ages, since we do not have exact age measurements. The grey area of the top panel corresponds to stellar spin-down without tidal interaction. Solid and dashed lines in the bottom panel correspond to cases with and without tidal interactions, respectively. The thickness of the solution curves (grey and colored) depends on the highest and lowest $\epsilon$ values in equation (1), which controls the sharpness of the transition in X-ray luminosity from the saturated (independent of stellar rotation) to the unsaturated (increases with stellar rotation) regimes.
@@
We find, therefore, that a close BD companion can significantly delay the spin-down of less magnetically active M dwarfs until tidal shredding. Such BD+M-dwarf systems can thus produce old M dwarf populations with short rotation periods ($<10$ days). In contrast, without a companion, the stellar wind removes enough angular momentum via magnetic braking to increase the period to hundreds of days, leading to fast and slow rotator branches at fixed age. Moreover, the rapid spin down after BD engulfment could explain the lack of observed M dwarfs with intermediate rotation periods for the subpopulation of M dwarfs.
@@
In addition to companion mass and orbital separation, the spin evolution is sensitive to the initial stellar rotation and magnetic field. Given that the combination of quantities needed for initial values in our model is not yet known for most M dwarfs, we identify values that produce model predictions that capture the observed gap best, while remaining within the observationally constrained range for each stellar property. For system 2 in Fig. 2 compared to Fig.1 (system 1), we change the initial rotation period to $0.3$ days and decrease the magnetic field to $50\rm G$. We keep the same companion mass range and start at $a_{\rm i}=2.9R_{\rm co}\approx 0.015\rm AU$ initial orbital separation. The spin evolves similarly to the previous case of Fig. 1, but slower. As such, we would not observe tidal shredding of heavier companions within the age of the universe. For companions more massive than $30M_{\rm J}$, M-dwarf spin-down is delayed and the star remains a fast rotator throughout its MS lifetime. To further illustrate the sensitivity of the spin evolution to the initial magnetic field strength and rotation period, we present complementary cases to the Figs. 1 and Fig. 2 in the supplementary material, varying the initial magnetic field strength and rotation period separately (see Figs. 6 and 7).
@@
For systems 1 and 2, we focused on M-dwarf+BD dynamical evolution for M dwarf initial rotation periods $P_{*}<1$day. In Fig. 3, we show results for system 3A which starts with an M-dwarf of $P_{*}=6$ days and $B_{\rm p,\star i}=200\rm G$, for the same companion BD mass range as in Fig. 2 at the orbital separation of $a_{\rm i}=0.5R_{\rm co}$. Similar to Fig. 1, Fig. 3 shows that the M dwarf initially spins down quickly from magnetic braking, reaching a quasi-equilibrium state $|\gamma_{\rm W}(t)|\simeq|\gamma_{\rm T}(t)|$, where $\gamma_{\rm W}(t)$ and $\gamma_{\rm T}(t)$ are the dimensionless stellar wind and tidal torque, respectively (equation 17). The companions migrate inward, spinning up the M dwarf but eventually engulfing. The M dwarf subsequently returns to magnetically mediated spin-down. As in Fig. 1, this can explain a rapid spin-down observed for M dwarfs (Newton et al., 2016). Here again, not many M-dwarfs would be expected along the steep rising curves on this plot, leading to a gap and fast and slow rotator populations indicated by the two schematically circled regions. Since initially tidal influence is not strong enough to significantly delay stellar spin-down or spin it up, the gap predicted from Fig. 3 would occur over a much shorter age range of only $7-10$ Gyr once companions get closer to the star, compared to that predicted in Fig. 1 and 2.
@@ 3.3 Spin evolution for partially convective stars
Jupiter mass companion planets are more common for K-dwarfs than for M-dwarfs. Here we consider system 4 of Table 1, a $0.6M_{\odot}$ mass star with a $10$ day rotation period, $5\rm G$ magnetic field, and a companion mass range $2M_{\rm J}\leq m\leq 16M_{\rm J}$ at an initial orbital separation $a_{i}=0.5R_{\rm co}$. Fig. 5 shows the resulting time evolution of the rotation period and X-ray activity. The plots exemplify that high-mass close companions can spin up a host star (top panel) and thereby increase X-ray activity (bottom panel). Compared to the M-Dwarf+BD systems of Fig. 2, this system produces more evenly distributed curves for a uniform companion mass distribution, so the gap discussed for those cases is not predicted here.
@@ 4.1 Substellar and Stellar companions of M Dwarfs
In general, observations of M dwarfs and their companions face several challenges due to their faintness, distances, and stellar variability. Combined with observational limitations, sample sizes and selection biases (Obermeier et al., 2016), these factors are likely to result in the underestimation of M dwarfs with companions (Cifuentes et al., 2025). Currently, the occurrence of known brown dwarfs at small separations ($\leq 1\rm{AU}$) around low-mass M dwarfs is approximately $1.3\%$ (Winters et al., 2019) and is even less for Jupiter mass planets (Endl et al., 2006). However, the stellar binary multiplicity rate for M dwarfs ranges from $28\%$ to $16\%$, depending on M dwarf mass (Winters et al., 2019). Recent observations (Cifuentes et al., 2025) suggest additional binary candidates which, if confirmed, could raise the multiplicity rate to $40.3\%$. This potentially significant increase suggests that more high-resolution observations are needed. In this study, we focus on M dwarfs with substellar companions for simplicity and to highlight the significance of tidal interactions in the context of spin evolution. It is a first step toward the future generalization of our model to stellar binary systems and the inclusion of mass accretion processes, given the high multiplicity rate of M dwarfs. Also, considering M dwarf systems with multiple small planetary companions, which are more frequent, could collectively produce enough tidal torque to significantly influence or even delay M dwarf spin-down.
@@ 4.2 Companion survival
Disk instability models suggest that Jupiter-mass planets can form rapidly around M dwarfs (Mercer and Stamatellos, 2020), so that the absence of Hot Jupiters around M dwarfs could be due to detection biases or post formation processes. High-energy radiation from active M dwarfs may lead to companion photoevaporation (Kurokawa and Nakamoto, 2014), while in less active M dwarf systems, such companions can migrate inward and eventually get engulfed by the host star (e.g. Siess and Livio, 1999; Nordhaus and Spiegel, 2013). In this study, we focus on less active M dwarfs, but according to our crude estimation based on energy-limited formalism from (Erkaev et al., 2007; Lopez and Fortney, 2013), even for highly active M dwarfs, companions with the masses studied here are unlikely to evaporate within 10 Gyr. Therefore, the planetary engulfment scenario seems a more probable outcome for our systems. Increased activity level and rotation of the star can be used as a proxy for tidal effects (Ilic et al., 2022), since inward migration and engulfment can lead to significant increase in stellar rotation (Siess and Livio, 1999; Carlberg et al., 2010; Oetjens et al., 2020) and activity (Siess and Livio, 1999; Kashyap et al., 2008; Ilić et al., 2023). Depending on stellar and companion properties, there may also be a temporary change in the surface abundance of lithium (Carlberg et al., 2010). (Cabezón et al., 2023) shows that BD engulfment by a solar-like star can cause lithium enhancement by a factor of 20-30, depending on the BD mass. Thus, in addition to stellar rotation and activity observations, high-precision abundance surveys are needed for companion host stars as indirect evidence of the recent substellar companion engulfment.

### v3

@@ 2.3 Combining magnetic braking and tides
$\times(1-2{\rm H}(\tilde{a}_{\rm crit}-\tilde{a}(t))).$ (19)
To account for the demise of the companion when it reaches the critical orbital separation $\tilde{a}_{\rm crit}$ defined as the larger of the Roche limit (Eggleton1983) or stellar radius, we include a multiplicative function $(1-2{\rm H}(\tilde{a}_{\rm crit}-\tilde{a}(t)))$ in the tidal torque term, where ${\rm H}(\tilde{a}_{\rm crit}-\tilde{a}(t))$ is the logistic function given by
@@
While the magnetic braking term always decreases the stellar spin, the influence of the tidal term depends on the orbital evolution of the secondary. If the secondary moves outward, then the orbital angular momentum will increase and tidal force will act to spin down the star. But if the secondary moves inward, reducing the orbital angular momentum, then tides increase the stellar spin. This evolution is thought to be much more efficient for cool stars with outer convection zones than for hot stars with outer radiative layers (Zahn2008).
Table 1: Fixed model parameters
Parameter and Definition Value Method References
[…]
$\color[rgb]{0,0,0}\lambda$ exponent of the power-law dependence of magnetic starspot area covering fraction on X-ray luminosity 1/3 Observation (NicholsFleming2020)
$\color[rgb]{0,0,0}q$ fraction of stellar mass anchored to the magnetic field 1 Theory Simple assumption
$\color[rgb]{0,0,0}\beta_{\star}$ radius of gyration 0.458, 0.359 111for $\color[rgb]{0,0,0}M_{\star}=0.2M_{\odot}$ & $\color[rgb]{0,0,0}M_{\star}=0.6M_{\odot}$. Simulation (Claret2023)
$\color[rgb]{0,0,0}k$ apsidal motion constant 0.155, 0.054 111for $\color[rgb]{0,0,0}M_{\star}=0.2M_{\odot}$ & $\color[rgb]{0,0,0}M_{\star}=0.6M_{\odot}$. Simulation (Claret2023)
[…]
@@ 3.1 Description of systems studied
To study M dwarf-BD and K dwarf-Jupiter mass companion systems we numerically solve the generalized system of equations (1), (4), (8), (9), (17) and (19) derived in section 2, which describe the stellar spin evolution when spin-down due to magnetic braking is coupled to tides for regime II. To distinguish between fully and partially convective cases, we vary stellar mass and take different values for certain fixed model parameters, while solving the same set of equations. A number of previous works indirectly infer that there may be difference in dynamo mechanisms between fully and partially convective stars (Vidotto2014; See2015; Folsom2018). Here we assume that dynamo-related differences may be higher-order corrections to the present effects. Incorporating this into our calculations is beyond the present scope. For all investigations below, we compare cases with both magnetic braking and tides to cases with just magnetic braking. For the latter, we solve equation (6) instead of equations (17) and (19). Our model parameters are given in Table 1 and the initial system properties for each case are given in Table 2. Using these values, we vary the companion mass and fix the initial value of the orbital separation, or vice versa. We define the initial orbital separation $a_{\rm i}$ as a fraction $p_{\rm f}$ of the corotation radius $a_{\rm i}=p_{\rm f}R_{\rm co}$, where the corotation radius is given by $R_{\rm co}=(G(M_{\star}+m)/\Omega^{2}_{\star})^{1/3}$.
@@
$P_{\star}/\rm Days$ 0.1 0.3 6 10 (popinchalk2021evaluating)
$B_{\rm p\star}/\rm G$ 100 50 200 5 (e.g., Vidotto2014; See2015; Kochukhov2020)
$\dot{M_{\star}}/\dot{M}_{\odot}$ 0.1 0.1 0.1 1 (Wood2021)
@@ 3.2 Spin evolution for fully convective stars
Figure 1: Evolution of rotation and orbital separation for System 1. We use a $0.2M_{\odot}$ mass M-dwarf with $0.1$ days initial period, $100\rm G$ magnetic field and companion BD at $a_{\rm i}=6R_{\rm co}$ initial orbital separation. Different colors represent varying BD companion masses between $15-75M_{\rm J}$. Top panel: spin evolution due only to magnetic braking (grey) and coupled to tides (colored). Cusps represent engulfment points. The thickness of the grey and colored curves is determined by the highest and lowest $\epsilon$ values (see text). Black dots represent fully convective M dwarfs with measured rotation periods and estimated ages from Galactic kinematics (irwin2011Angular). The vertical grey dashed lines separate M-dwarfs in the thin, intermediate and thick disk populations with corresponding age ranges of 0.530.5-3 Gyr, 373-7 Gyr and 7137-13 Gyr. The x-axis and y-axis are normalized to 1 Gyr and a 1-day rotation period, respectively. Bottom panel: orbital separation for solutions without tides (dashed) and coupled to tides (solid) analgous to the top panel. The x-axis is normalized to 1 Gyr, and the y-axis is normalized to the stellar radius $R_{\star}$.
@@
The top and bottom panels of Fig. 1 show the time evolution of the rotation period and orbital separation for system 1 of Table 2. For initial values we take a rotation period of $P_{\star}=0.1$ days, a magnetic field $B_{\rm p,\star i}=100\rm G$ and orbital separation $a_{\rm i}=6R_{\rm co}\approx 0.015\rm AU$. We consider cases of companion BD masses between $15-75M_{\rm J}$, each represented by a different color. Mass accretion processes are not considered in the current study, and may be important for large companion masses. We take $75M_{\rm J}$ BD dwarf as our largest companion mass case for visualization and to serve as a conceptual bridge towards future work involving stellar binaries and mass transfer. Black dots are data points from (irwin2011Angular) with measured rotation periods, and estimated ages from galactic kinematics for 41 field M-dwarfs. This data set typifies the observed trend, and provides a simple comparison to our results, given the availability of both rotation periods and ages in the sample. In Fig. 1, the vertical grey dashed lines represent boundaries between Galactic thin, intermediate, and thick disk M-dwarf populations with assigned ages 0.530.5-3 Gyr, 373-7 Gyr, and 7137-13 Gyr, respectively. The data points in Fig. 1 use the average value of these disk sub-population ages, since we do not have exact age measurements. The grey area of the top panel corresponds to stellar spin-down without tidal interaction. Solid and dashed lines in the bottom panel correspond to cases with and without tidal interactions, respectively. The thickness of the solution curves (grey and colored) depends on the highest and lowest $\epsilon$ values in equation (1), which controls the sharpness of the transition in X-ray luminosity from the saturated (independent of stellar rotation) to the unsaturated (increases with stellar rotation) regimes.
@@
We find, therefore, that a close BD companion can significantly delay the spin-down of less magnetically active M dwarfs until tidal shredding. Such BD+M-dwarf systems can thus produce old M dwarf populations with short rotation periods (<10<10 days). In contrast, without a companion, the stellar wind removes enough angular momentum via magnetic braking to increase the period to hundreds of days, leading to fast and slow rotator branches at fixed age. Moreover, the rapid spin down after BD engulfment could explain the lack of observed M dwarfs with intermediate rotation periods for the subpopulation of M dwarfs.
@@
In addition to companion mass and orbital separation, the spin evolution is sensitive to the initial stellar rotation and magnetic field. Given that the combination of quantities needed for initial values in our model is not yet known for most M dwarfs, we identify values that produce model predictions that capture the observed gap best, while remaining within the observationally constrained range for each stellar property. For system 2 in Fig. 2 compared to Fig.1 (system 1), we change the initial rotation period to $0.3$ days and decrease the magnetic field to $50\rm G$. We keep the same companion mass range and start at $a_{\rm i}=2.9R_{\rm co}\approx 0.015\rm AU$ initial orbital separation. The spin evolves similarly to the previous case of Fig. 1, but slower. As such, we would not observe tidal shredding of heavier companions within the age of the universe. For companions more massive than $30M_{\rm J}$, M-dwarf spin-down is delayed and the star remains a fast rotator throughout its MS lifetime. To further illustrate the sensitivity of the spin evolution to the initial magnetic field strength and rotation period, we present complementary cases to the Figs. 1 and Fig. 2 in the supplementary material, varying the initial magnetic field strength and rotation period separately (see Figs. 7 and 8).
@@
For systems 1 and 2, we focused on M-dwarf+BD dynamical evolution for M dwarf initial rotation periods $P_{\star}<1$day. In Fig. 3, we show results for system 3A which starts with an M-dwarf of $P_{\star}=6$ days and $B_{\rm p,\star i}=200\rm G$, for the same companion BD mass range as in Fig. 2 at the orbital separation of $a_{\rm i}=0.5R_{\rm co}$. Similar to Fig. 1, Fig. 3 shows that the M dwarf initially spins down quickly from magnetic braking, reaching a quasi-equilibrium state $|\gamma_{\rm W}(t)|\simeq|\gamma_{\rm T}(t)|$, where $\gamma_{\rm W}(t)$ and $\gamma_{\rm T}(t)$ are the dimensionless stellar wind and tidal torque, respectively (equation 19). The companions migrate inward, spinning up the M dwarf but eventually engulfing. The M dwarf subsequently returns to magnetically mediated spin-down. As in Fig. 1, this can explain a rapid spin-down observed for M dwarfs (Newton2016). Here again, not many M-dwarfs would be expected along the steep rising curves on this plot, leading to a gap and fast and slow rotator populations indicated by the two schematically circled regions. Since initially tidal influence is not strong enough to significantly delay stellar spin-down or spin it up, the gap predicted from Fig. 3 would occur over a much shorter age range of only 7107-10 Gyr once companions get closer to the star, compared to that predicted in Figs. 1 and 2.
@@
In our model we use a simpler prescription for tidal interactions and do not take into account structural evolution of the host star or dynamical tides as compared to previous work (Gallet2013; Gallet+Bouvier2015; Amard2016; Garraffo2018; benbakoura2019evolution; ahuir2021dynamical; Pezzotti2025). Although we have focused on lower mass star systems, our simpler self-contained analytic approach, when applied to solar-like stars, captures similar trends for the influence of tides. See for example Fig. 10 from ahuir2021dynamical and our Fig. 1, but note the difference in system parameters. The ultimate limitations of accuracy from different assumptions and ingredients that go into these different frameworks makes the present simpler approach complementarily useful. A more detailed comparison between these different approaches across parameter regimes is beyond the present scope.
@@ 3.3 Spin evolution for partially convective stars
Jupiter mass companion planets are more common for K-dwarfs than for M-dwarfs. Here we consider system 4 of Table 2, a $0.6M_{\odot}$ mass star with a $10$ day rotation period, $5\rm G$ magnetic field, and a companion mass range $2M_{\rm J}\leq m\leq 16M_{\rm J}$ at an initial orbital separation $a_{i}=0.5R_{\rm co}$. Fig. 5 shows the resulting time evolution of the rotation period and X-ray activity. The plots exemplify that high-mass close companions can spin up a host star (top panel) and thereby increase X-ray activity (bottom panel). Our results agree with previous studies (e.g ilic2022tidal; Pezzotti2025), which show that massive close-in planets can increase X-ray activity and shorten rotation periods. Compared to the M-Dwarf+BD systems of Fig. 2, this system produces more evenly distributed curves for a uniform companion mass distribution, so the gap discussed for those cases is not predicted here.
@@ 4.1 Substellar and Stellar companions of M Dwarfs
In general, observations of M dwarfs and their companions face several challenges due to their faintness, distances, and stellar variability. Combined with observational limitations, sample sizes and selection biases (Obermeier2016), these factors are likely to result in the underestimation of M dwarfs with companions (Cifuentes2025carmenes). Currently, the occurrence of known brown dwarfs at small separations ($\leq 1\rm{AU}$) around low-mass M dwarfs is approximately $1.3\%$ (Winters+2019) and is even less for Jupiter mass planets (Endl2006). However, the stellar binary multiplicity rate for M dwarfs ranges from $28\%$ to $16\%$, depending on M dwarf mass (Winters+2019). Recent observations (Cifuentes2025carmenes) suggest additional binary candidates which, if confirmed, could raise the multiplicity rate to $40.3\%$. This potentially significant increase suggests that more high-resolution observations are needed. In this study, we focus on M dwarfs with substellar companions for simplicity and to highlight the significance of tidal interactions in the context of spin evolution. It is a first step toward the future generalization of our model to stellar binary systems and the inclusion of mass accretion processes, given the high multiplicity rate of M dwarfs. Also, considering M dwarf systems with multiple small planetary companions, which are more frequent, could collectively produce enough tidal torque to significantly influence or even delay M dwarf spin-down.
@@ 4.2 Companion survival
Disk instability models suggest that Jupiter-mass planets can form rapidly around M dwarfs (Mercer2020A), so that the absence of Hot Jupiters around M dwarfs could be due to detection biases or post formation processes. High-energy radiation from active M dwarfs may lead to companion photoevaporation (Kurokawa2014), while in less active M dwarf systems, such companions can migrate inward and eventually get engulfed by the host star (e.g. Siess1999; Nordhaus2013). In this study, we focus on less active M dwarfs, but according to our crude estimation based on energy-limited formalism from (Erkaev2007; Lopez2013), even for highly active M dwarfs, companions with the masses studied here are unlikely to evaporate within 10 Gyr. Therefore, the planetary engulfment scenario seems a more probable outcome for our systems. Increased activity level and rotation of the star can be used as a proxy for tidal effects (Ilic2022), since inward migration and engulfment can lead to significant increase in stellar rotation (Siess1999; Carlberg2010; Oetjens2020) and activity (Siess1999; Kashyap2008; Ilic2023). Depending on stellar and companion properties, there may also be a temporary change in the surface abundance of lithium (Carlberg2010). (Cabezon2023) shows that BD engulfment by a solar-like star can cause lithium enhancement by a factor of 20-30, depending on the BD mass. Thus, in addition to stellar rotation and activity observations, high-precision abundance surveys are needed for companion host stars as indirect evidence of the recent substellar companion engulfment.